# Experience-Driven Exploration for Efficient API-Free AI Agents

### v2

@@ 3.3 Reasoning and Decision-Making Module
$\mathbb{S}_{C}(v_{i})=\{\sigma_{k}\mid\sigma_{k}\in\mathbb{S}\ \land\ f_{VLM}(prompt^{invoke},X_{i},\mathbb{S})=\sigma_{k}\},$ (5)
where $\mathbb{S}_{C}(v_{i})$ denotes the candidate skills retrieved by the VLM based on the visual observation $X_{i}$. To avoid committing greedily in a stochastic and partially observable environment, the agent evaluates the candidate skills $\mathbb{S}_{C}(v_{i})$ using an approach inspired by the Upper Confidence bound for Trees (UCT) algorithm Couëtoux et al. (2011), a core component of MCTS. For each candidate skill $\sigma_{k}\in\mathbb{S}_{C}(v_{i})$, the agent calculates a potential utility score which balances exploitation of known, high-value skills with the exploration of less-tried options. This score $\eta_{\sigma_{k}}$ is defined as:
@@
where the $\phi(\sigma_{k})$ is the estimated fitness of skill $\sigma_{k}$. The second term is the exploration bonus, where the $n_{k}$ is the execution count for skill $\sigma_{k}$, the $N_{i}$ is the total number of selections $\mathbb{S}_{C}(v_{i})$, and the $C$ is a constant controlling the exploration-exploitation trade-off. $P(\sigma_{k})$ is a penalty term that dynamically reduces the utility of skills whose prerequisite objects are not currently available, promoting the selection of fully completable actions. Finally, instead of deterministically picking the skill with the highest utility, the agent converts these scores into a probability distribution using a temperature-scaled softmax function. The probability of selecting skill $\sigma_{k}$ is given by:
@@
The skill to be executed is then stochastically sampled from this distribution. The temperature parameter $\tau$ dynamically adjusts the randomness of the selection. This method ensures a robust balance between exploiting proven strategies and exploring new possibilities. If the candidate set $\mathbb{S}_{C}(v_{i})$ is empty, the agent defaults to skill augmentation in the current context.
Skill Augmentation & Refinement. In open-ended environments, most atomic actions $a\in\mathcal{A}$ are inherently task-irrelevant and semantically meaningless in the absence of predefined APIs and priors. Therefore, discovering useful skills $\sigma=(a_{1},\ldots,a_{k})$ necessitates a trial-and-reasoning process: the agent explores diverse combinations of atomic actions and reflects on their outcomes to identify meaningful behaviors. To cope with the intractably large action space, besides leveraging SAM to identify and segment UI elements and potential interaction targets, we also construct skills through incremental increases in sequence length $k$. The process begins by evaluating single-step actions ($k=1$), then progressively expands to longer sequences ($k=2,3,\ldots$), while aggressively pruning unproductive branches based on environmental feedback. Only those sequences that yield recognizable and valuable effects, e.g., GUI transitions, measurable task progress, or meaningful state changes in SA-KG, are retained as useful skills. Each new candidate skill $\sigma_{k}$ is formed by appending a new atomic action $a_{k}$ to a validated shorter skill $\sigma_{k-1}$. This iterative process of expansion and pruning ensures that the skill library maintains only potentially useful routines, allowing complex behaviors to emerge compositionally from primitive actions. When the set of practicable skills grows sparse, the agent triggers the VLM to generate a semantically refined variant as follows:
$\sigma^{\prime}=f_{VLM}(prompt^{refine},(X_{i},\sigma,\mathcal{T})).$ (8)
@@
The dual process, i.e., combining population-based pruning with LLM-guided rewriting, ensures that $\mathbb{S}$ undergoes evolution towards a more reusable, and semantically coherent skill repertoire.
@@
$\mathcal{R}_{total}=\mathcal{R}_{progress}+\mathcal{R}_{semantics}+\mathcal{R}_{state}+\mathcal{R}_{novel},$ (9)
where $\mathcal{R}_{progress}$ evaluates the progress of the game via VLM to encourage the skill to lead to specific progress in the game, $\mathcal{R}_{semantics}$ measures the consistency via VLM between the predicted high-level effects of the skill and the actual results observed in the environment, $\mathcal{R}_{state}$ encourages the agent to execute skills that lead to states with greater future potential, $\mathcal{R}_{novel}$ serves as an intrinsic curiosity drive, directly rewarding the agent to explore and expand its knowledge.
To compute $\mathcal{R}_{state}$, we define a static skill potential $\mathcal{V}_{S}(v_{i})$ as the sum of the weights of all outgoing skill edges. Then, the state-value reward for a transition from $v_{i}$ to $v_{j}$ is defined as the improvement in the agent’s estimated long-term potential:
$\mathcal{R}_{state}=\mathcal{V}_{S}(v_{j})-\mathcal{V}_{S}(v_{i})=\sum_{e\in\mathcal{E}^{\sigma}(v_{j})}w^{\sigma}(e)-\sum_{e\in\mathcal{E}^{\sigma}(v_{i})}w^{\sigma}(e).$ (10)
@@
[…]
@@
(Floors) Scores Responsive Rate Costs (Turns) Researched Responsive Rate Costs
GPT-4o* 8 46 0.51 1.0 13 1 0.57 0.9
Claude3.7* 1 5 0.80 1.2 17 3 0.93 1.2
UITARS-1.5* 1 5 0.53 0.7 10 1 0.93 0.1
GPT-4o 1 5 0.71 1.2 6 0 0.36 0.7
Claude3.7 1 5 0.26 1.0 0 0 0.15 1.0
UITARS-1.5 1 5 0.81 0.1 0 0 0.62 0.1
BottomUp 13 81 0.99 2.4 50 8 0.92 3.0
KG-Agent 16 112 0.99 2.1 115 15 0.94 2.8

### v3

@@ III-C Reasoning and Decision-Making Module
$\mathbb{S}_{C}(v_{i})=\{\sigma_{k}\mid\sigma_{k}\in\mathbb{S}\ \land\ f_{*}(\mathbb{P}^{invoke},X_{i},\mathbb{S})=\sigma_{k}\},$ (5)
where $\mathbb{S}_{C}(v_{i})$ denotes the candidate skills retrieved by the VLM based on the visual observation $X_{i}$. To avoid committing greedily in a stochastic and partially observable environment, the agent evaluates the candidate skills $\mathbb{S}_{C}(v_{i})$ using an approach inspired by the Upper Confidence bound for Trees (UCT) algorithm [37], a core component of MCTS. For each candidate skill $\sigma_{k}\in\mathbb{S}_{C}(v_{i})$, the agent calculates a potential utility score which balances exploitation of known, high-value skills with the exploration of less-tried options. This Upper Confidence Bound (UCB) score $\eta_{\sigma_{k}}$ is defined as:
@@
where $\phi(\sigma_{k})$ quantifies the fitness of skill $\sigma_{k}$. Initially set to 0, its value is updated based on two factors: detectable visual state transitions following execution, and the consistency of the skill’s semantic description with the outcomes observed across before-and-after frames. The second term is the exploration bonus, where the $n_{k}$ is the execution count for skill $\sigma_{k}$, the $N_{i}$ is the total number of selections $\mathbb{S}_{C}(v_{i})$, and the $C$ is a constant controlling the exploration-exploitation trade-off. $P(\sigma_{k})$ is a penalty term that dynamically reduces the utility of skills whose prerequisite objects are not currently available, promoting the selection of fully completable actions. Finally, instead of deterministically picking the skill with the highest utility, the agent converts these scores into a probability distribution using a temperature-scaled softmax function. The probability of selecting skill $\sigma_{k}$ is given by:
@@
The skill to be executed is then stochastically sampled from this distribution. The temperature $\tau$ dynamically adjusts the randomness of the selection. This method ensures a robust balance between exploiting proven strategies and exploring new possibilities. If the candidate set $\mathbb{S}_{C}(v_{i})$ is empty, the agent defaults to skill augmentation in the current context.
Skill Augmentation & Refinement. In open-ended environments, where predefined APIs and priors are absent, most atomic actions $a\in\mathcal{A}$ are inherently task-irrelevant and semantically ambiguous. Consequently, discovering useful skills $\sigma=(a_{1},\ldots,a_{k})$ necessitates a structured trial-and-reasoning process, in which the agent explores diverse action combinations and evaluates their outcomes to identify meaningful behaviors. To manage the intractably large action space, we adopt the strategy by leveraging SAM to identify and segment UI elements, as well as constructing skills through incremental increases in sequence length $k$. Notably, our skill augmentation approach is driven by successful atomic action creation and validation rather than exhaustive combinatorial search. We partition the atomic action set $\mathcal{A}$ into three ordered subsets, i.e., $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}$, to optimize the search space and accelerate exploration. The agent begins with $\mathcal{A}_{1}$, which contains previously validated single-step skills, then progresses to $\mathcal{A}_{2}$, comprising actions on recognized UI objects that prioritize semantically meaningful interactions, and finally considers $\mathcal{A}_{3}$, which includes residual actions such as clicks on unrecognized or background regions. The skill construction process starts with single-step actions ($k=1$) and incrementally expands to longer sequences ($k=2,3,\ldots$). Each candidate skill $\sigma_{k}$ is formed by appending a new atomic action $a_{k}$ to a validated shorter skill $\sigma_{k-1}$. Skill expansion terminates as soon as an action sequence produces any recognizable and valuable effect, e.g., GUI transitions, measurable task progress, or meaningful state changes in the SAG—at which point the skill is annotated functionally and stored in memory.
To maintain an efficient and scalable skill library, we implement a dynamic skill pruning mechanism. MCTS is employed to ensure all skills receive sufficient testing opportunities. During skill execution, the agent records trajectory data and computes UCB scores according to Eq. 6. Skills that exceed the average visitation count and consistently demonstrate the lowest UCB scores are pruned from memory. This iterative process of expansion and pruning ensures that the skill library retains only high-utility behaviors, enabling complex, adaptive strategies to emerge compositionally from primitive actions. Furthermore, to maintain a compact and non-redundant skill set, the agent periodically triggers the VLM to consolidate the skill repertoire by clustering and merging functionally equivalent skills. This process is defined as:
$\sigma^{\prime}=f_{*}(\mathbb{P}^{refine},(X_{i},\sigma,\mathcal{T})).$ (8)
@@
The dual process, i.e., combining population-based pruning with LLM-guided skill consolidation, ensures that $\mathbb{S}$ undergoes evolution towards a more reusable and semantically coherent state by eliminating redundancy and promoting general-purpose skills.
Input: Environment $(\mathcal{X},\mathcal{A},\mathcal{T},\mathcal{R})$, VLM $f_{*}$, Skill library $\mathbb{S}=\emptyset$, and SAG $\mathcal{G}(\mathcal{V,E})=\emptyset$.
1 for Each Agent do
2 while Episode not terminated do
3      Observe screen $X_{i}$ and extract feature $x_{i}$;
4      Generate node $v_{i}$ and similarity edges $e^{sim}$;
5      Identify Neighborhood of Experience by Eq. 2;
6      Gather High-Quality Candidate Skills by Eq. 3;
7    if $\mathbb{S}_{HQ}(v_{i})\neq\emptyset$ then
8       for $attempt=1$ to max_attemptsmax\_attempts do
9            Sample high-quality skill $\sigma_{k}$ by Eq. 4;
10            Compute $\mathcal{R}_{total}$ by Eq. 9;
11            Generate skill edges $w^{\sigma_{k}}_{i,j}$ by Eq. 1;
12          if $\mathcal{R}_{total}(v_{i},v_{j})>threshold$ then
13             Break;
14
15
16
17    else
18         Sample skill $\sigma_{l}$ from $\mathbb{S}_{C}(v_{i})$ by Eq. 7;
19       if $\mathbb{S}_{C}(v_{i})=\emptyset$ then
20          for $k=1$ to $k_{max}$ do
21               Gen. $a\leftarrow f_{*}(\mathbb{P}^{augment},X_{i},\mathbb{O})$;
22             if $a$ is recognizablerecognizable then
23                  Update Skill library $\mathbb{S}$;
24                break;
25
26
27
28       else
29            Select skill cluster by $f_{*}(\mathbb{P}^{invoke},X,\mathbb{S})$;
30            Eval. the skill cluster with MCTS;
31            Execute the best skill $\sigma$ from MCTS;
32
33        Observe $\mathcal{T}$ & compute $\mathcal{R}_{total}$ by Eq. 9;
34       if $\mathcal{R}_{total}(v_{i},v_{j})<threshold$ then
35            Remove $\sigma$ from $\mathbb{S}$;
36
37       if skill_count<thresholdskill\_count<threshold then
38            Refine skill: $\sigma^{\prime}\leftarrow f_{*}(\mathbb{P}^{refine},X,\sigma,\mathcal{T})$;
39            Replace $\sigma$ with $\sigma^{\prime}$ if improvement;
40
41
42
43
Algorithm 1 Skill Evolution of SAG-Agent.
@@ 39            Replace $\sigma$ with $\sigma^{\prime}$ if improvement;
$\mathcal{R}_{total}=\mathcal{R}_{progress}+\mathcal{R}_{semantic}+\mathcal{R}_{state}+\mathcal{R}_{novel},$ (9)
where $\mathcal{R}_{progress}$ evaluates the progress of the game via VLM to encourage the skill to lead to specific progress in the game, $\mathcal{R}_{semantic}$ measures the consistency via VLM between the predicted high-level effects of the skill and the actual results observed in the environment, $\mathcal{R}_{state}$ encourages the agent to execute skills that lead to states with greater future potential, $\mathcal{R}_{novel}$ serves as an intrinsic curiosity drive, directly rewarding the agent to explore and expand its knowledge.
To compute $\mathcal{R}_{state}$, we define a static skill potential as the sum of the weights of all outgoing skill edges. Then, the state-value reward for a transition from $v_{i}$ to $v_{j}$ is defined as the improvement in the agent’s estimated long-term potential:
$\mathcal{R}_{state}=\sum_{e\in\mathcal{E}^{\sigma}(v_{j})}w^{\sigma}(e)-\sum_{e\in\mathcal{E}^{\sigma}(v_{i})}w^{\sigma}(e).$ (10)
@@
III-D Skill Evolution of SAG-Agent
As shown in Algorithm 1, a core innovation of our framework is the derivation of the $\mathcal{N}_{\mathcal{E}}(v_{i})$, which enables the agent to leverage past experience by aggregating a set of high-quality candidate skills $\mathbb{S}{HQ}(v_{i})$ from the SAG, specifically, skills proven effective in analogous states. The agent then executes these skills, with the resulting reward feedback updating the corresponding skill edges in the graph. The loop terminates early upon achieving a sufficiently high reward, ensuring efficient skill reuse. When no such skills are available, the agent enters a trial-and-reasoning phase guided by a UCT-style selector within MCTS. This selector uses a fixed exploration constant $C=0.5$ and a base exploration utility threshold of $0.1$, while dynamically scheduling the temperature $\tau$ via a decay function based on the selection count of the relevant skill cluster to shift gradually from exploration to exploitation. Beyond standard UCT, the selector incorporates specialized mechanisms: a pre-selection filter excludes actions requiring unavailable or suspended objects; an action completeness penalty reduces fitness scores based on missing prerequisites; and an exploration dominance trigger activates pure exploration when the probability for exploratory actions exceeds $0.9$. These thresholds are coupled with numerical stability measures, e.g., max normalization and probability clipping in the softmax function, to ensure robust action selection in complex, partially observable environments. This integrated approach allows the SAG-Agent to dynamically balance the exploitation of proven strategies with the exploration of novel solutions, while maintaining an evolving knowledge structure that captures both state relationships and skill effectiveness.
Figure 3: Prompt design for VLM-based reasoning stages: skill invocation, augmentation, refinement, and evaluation.
TABLE I: Performance comparison across Slay the Spire and Civilization V. To validate performance, each agent is evaluated over three independent episodes per environment, as well as the mean and standard deviation of the final results are reported. Methods with * indicate prior-assisted setting, while those without * indicate zero-prior. NA denotes not applicable.
@@ III-D Skill Evolution of SAG-Agent
GPT-4o* $8.33\pm 0.58$ $48.73\pm 2.43$ $0.49\pm 0.02$ $1.01\pm 0.03$ $14.67\pm 2.08$ $1.00\pm 0.00$ $0.57\pm 0.01$ $0.91\pm 0.02$
Claude3.7* $1.00\pm 0.00$ $5.00\pm 0.00$ $0.79\pm 0.01$ $1.22\pm 0.04$ $19.33\pm 2.52$ $3.33\pm 0.58$ $0.92\pm 0.01$ $1.18\pm 0.09$
UITARS-1.5* $1.00\pm 0.00$ $5.00\pm 0.00$ $0.54\pm 0.01$ $0.49\pm 0.24$ $11.33\pm 1.15$ $1.33\pm 0.58$ $0.92\pm 0.02$ $0.12\pm 0.01$
CRADLE* $1.00\pm 0.00$ $5.00\pm 0.00$ NA $4.89\pm 0.23$ $\pm 0.00 0.00±0.00$$\pm 0.00$ NA $4.23\pm 0.28$
GPT-4o $1.00\pm 0.00$ $5.00\pm 0.00$ $0.72\pm 0.03$ $1.27\pm 0.09$ $6.33\pm 0.58$ $0.00\pm 0.00$ $0.37\pm 0.04$ $0.76\pm 0.03$
Claude3.7 $1.00\pm 0.00$ $5.00\pm 0.00$ $0.28\pm 0.01$ $1.08\pm 0.06$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $0.15\pm 0.3$ $1.22\pm 0.13$
UITARS-1.5 $1.00\pm 0.00$ $5.00\pm 0.00$ $82\pm 0.01 0.0$$9\pm 0.01 0.0$$\pm 0.00 0.00±0$$0\pm 0.00$ $0.63\pm 0.01$ $0.10\pm 0.01$
BottomUp $14.33\pm 1.15$ $91.33\pm 9.61$ $0.98\pm 0.01$ $2.68\pm 0.29$ $59.67\pm 8.74$ $8.66\pm 0.58$ $0.92\pm 0.01$ $3.55\pm 0.52$
SAG-Agent $15.67\pm 0.47$ $104.00\pm 5.89$ $0.98\pm 0.01$ $2.15\pm 0.04$ $107.33\pm 6.55$ $13.33\pm 1.25$ $0.93\pm 0.01$ $3.03\pm 0.17$
III-E Prompt Engineering for Agent Reasoning
Effective reasoning in VLMs relies on meticulously designed prompts to guide complex visual understanding and decision-making processes [38]. In SAG-Agent, we employ a structured VLM-based reasoning cycle, i.e., skill invocation, augmentation, refinement, and evaluation. As shown in Fig. 3, each stage is supported by a dedicated prompt that collectively enables intelligent interaction in dynamic GUI environments without reliance on pre-defined APIs. Specifically, the prompt for skill invocation selects the most suitable executable action from procedural memory when symbolic reasoning over the SAG fails. The prompt for skill augmentation complements invocation by deriving new high-level actions from low-level UI operations, thereby expanding the agent’s skill repertoire based on observed screen changes. The prompt for skill refinement consists of two sub-prompts: one groups functionally similar new actions, and the other merges them with existing skill clusters to maintain a compact and non-redundant skill library. Finally, the prompt for skill evaluation assesses the effectiveness of executed skills by comparing expected outcomes against actual screen changes, providing essential feedback for progressive and semantic rewards. Together, these prompts form a closed-loop learning architecture: invocation drives action, augmentation acquires knowledge, refinement consolidates understanding, and evaluation validates progress.
IV Experiments
IV-A Environment and Experimental Setup
Environment. We conduct experiments in two complex, open-ended games established by the Bottom-up Agent [13]: Slay the Spire (Base difficulty: Ascension 0, Character: Ironclad), a roguelike deck-builder, and Civilization V (Civilization: Russia, Map: Earth, Size: Standard, Difficulty: Prince, Pace: Standard), a hallmark 4X strategy game. These environments serve as ideal testbeds as they demand both precise, low-level execution and sophisticated, long-horizon planning. Their turn-based nature is also well-suited to the current latency characteristics of LLM-based reasoning. An episode is terminated if the agent fails a task-specific condition (character defeat in Slay the Spire or nation defeat in Civilization V) or if it reaches a maximum limit of 500 steps.
[…]
Baselines. We evaluate SAG-Agent against a comprehensive suite of representative baselines. This includes leading proprietary models (GPT-4o, Claude 3.7), a strong open-source GUI agent (UI-TARS-1.5 [1]), and the two most relevant API-free agents (CRADLE [12] and the Bottom-up Agent [13]). This selection is motivated by the limitations of many existing frameworks, which are often incompatible with the API-free setting or require task-specific priors (e.g., predefined subgoals). For the CRADLE baseline, which natively depends on a library of atomic skills, we adapted its core prompts for our tasks. To ensure a fair comparison, we converted the skills learned by SAG-Agent into a compatible format that could be utilized by the CRADLE framework, thereby providing it with the necessary prerequisite capabilities.
Implementation Details. We use GPT-4o for semantic reasoning and a pre-trained CLIP ViT-B/32 for visual encoding. All hyperparameters remain consistent across all experiments to ensure fair and reproducible comparisons. For the SAG, the state merging threshold $\theta_{\text{merge}}$ is set to 0.95, and the similarity edge threshold $\theta_{\text{simi}}$ to 0.88. For skill edge weight computation (Eq. 1), we use a balancing factor $\alpha=0.7$ and fitness sensitivity constant $C_{0}=5.0$. The UCT-based skill selection (Eq. 6) uses an exploration constant $C_{1}=5.0$. Critically, the exact same agent implementation, including the identical hyperparameter configuration and prompt design, is deployed in both Slay the Spire and Civilization V. This highlights the domain-general nature of our framework, which operates effectively without environment-specific tuning.
IV-B Comparative Results
As shown in Table I, the SAG-Agent demonstrates statistically significant and consistent superiority over all baseline methods across two distinct open-ended game environments, including its zero-shot transfer to Slay the Spire under identical architecture, prompts, and hyperparameters as used in Civilization V, highlighting the framework’s inherent adaptability and robustness without any domain-specific adjustments. Our agent achieves the highest progression metrics and maximum in-game scores while maintaining exceptional execution reliability, with performance consistency across different game genres further underscoring the robustness of our approach. The contrast with CRADLE is particularly illuminating: while CRADLE relies on extensive prior knowledge encoded in hand-designed prompts and atomic skills, it achieves only minimal progression in both games. SAG-Agent’s substantial advantage without any game-specific prior knowledge validates the effectiveness of our zero-prior learning paradigm. Furthermore, our method demonstrates remarkable consistency, achieving top performance across all metrics in both games, unlike baselines that exhibit volatile performance between environments. Notably, SAG-Agent accomplishes this superior performance with reduced computational overhead, evidenced by lower token costs per 100 steps compared to the BottomUp Agent. These results confirm that our agent not only inherits BottomUp’s execution reliability but significantly enhances strategic decision-making through our skill-aligned knowledge graph, ultimately affirming SAG-Agent’s effectiveness in mastering complex, open-ended environments through autonomous learning and adaptation.
TABLE II: Ablation study of core components of SAG-Agent in Civilization V, with each run constrained to 100 steps for clarity. The “w/o Simi.” means that each state forms an isolated node, with no merging or similarity-based connections permitted. $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ denote $\mathcal{R}_{progress}$ and $\mathcal{R}_{semantic}$, respectively. $\mathcal{R}_{3}$ and $\mathcal{R}_{4}$ denote $\mathcal{R}_{state}$ and $\mathcal{R}_{novel}$, respectively.
Setting Library SAG Information Civilization V
[…]
Full Model 76 26 37 226 65 10 0.89 3.0
w/o Simi. 30 672 62 0 15 1 0.64 3.5
w/o $\mathcal{R}_{4}$ 48 11 19 74 23 2 0.81 3.1
w/o $\mathcal{R}_{3}$ 57 20 33 110 39 4 0.87 3.2
w/o $\mathcal{R}_{3}$ & $\mathcal{R}_{4}$ 64 18 39 96 35 3 0.79 3.3
w/o $\mathcal{R}_{1}$ & $\mathcal{R}_{2}$ 33 9 14 42 14 1 0.59 2.1
TABLE III: Sensitivity analysis of key hyperparameters in Slay the Spire, including the number of skills sampled per execution attempt $M$, the coefficient balancing state-value reward $\mathcal{R}_{state}$ ($\mathcal{R}_{3}$) and novelty reward $\mathcal{R}_{novel}$ ($\mathcal{R}_{3}$), and the cosine-similarity thresholds for node merging and similarity edge creation ($\theta_{merge}$, $\theta_{simi}$). Time is the average time in seconds required to run one step under each configuration.
Configuration SAG Information Slay the Spire
[…]
$M=5$ 29 54 354 16 112 158.34s
$M=3$ 25 38 124 16 98 244.63s
$M=10$ 22 34 116 11 63 192.41s
$\mathcal{R}_{3}:\mathcal{R}_{4}=1:1$ 29 54 354 16 112 158.34s
$\mathcal{R}_{3}:\mathcal{R}_{4}=1:2$ 33 32 220 16 102 153.26s
$\mathcal{R}_{3}:\mathcal{R}_{4}=2:1$ 20 26 114 13 87 145.92s
(0.95, 0.88) 29 54 354 16 112 158.34s
(0.99, 0.90) 155 190 14520 12 70 392.55s
(0.85, 0.80) 4 9 4 4 26 179.40s
IV-C Ablation Studies
As shown in Table II, the full KG‑Agent achieves the strongest performance, while removing core components leads to marked degradation. Specifically, disabling similarity-based merging (“w/o Simi.”) severely impairs long‑term progression and strategic coherence. This drop stems from a deeper architectural inconsistency: although state merging was disabled, the graph‑based novelty reward $R_{novel}$ ($R_{4}$) and state‑transition reward $R_{state}$ ($R_{3}$) remained active. The resulting mismatch caused $R_{novel}$ ($R_{4}$) to continuously reward the agent for registering each new state as an isolated node, promoting graph expansion over semantic exploration, while $R_{state}$ ($R_{3}$) operated on a fragmented graph with disconnected nodes, yielding unstable and localised value estimates. Consequently, the VLM was misled by pathological reward signals, prioritising node proliferation over meaningful progress, as reflected in the inflated node count (672 nodes vs. 26 in the full model). Similarly, ablating $\mathcal{R}{progress}$ ($R_{1}$) and $\mathcal{R}{semantic}$ ($R_{2}$) reduces the total reward available for skill evaluation, which directly weakens skill fitness $\phi(\sigma_{k})$ and lowers the corresponding UCB scores during MCTS. This impairs both skill selection and pruning, resulting in notably poor outcomes across library size, graph connectivity, and in‑game progression. Together, these ablation studies confirm that our complete framework, integrating similarity‑based abstraction with a balanced multi‑reward mechanism, enables consistent state representation and stable value estimation, thereby providing reliable learning signals for effective long‑horizon planning.
TABLE IV: Evolution of skill library and SAG, as well as game performance over training rounds in Civilization V. Each round builds on previous memory with 100 steps.
Round Skill Library Information SAG Information Civilization V
[…]
Round 0 76 +77 -1 26 37 226 65 10 0.89 3.0
Round 1 106 +33 -3 50 84 782 98 12 0.92 2.8
Round 2 111 +7 -2 53 90 856 113 13 0.93 2.7
Round 3 114 +4 -1 55 90 838 115 13 0.94 2.7